Session: May 1997 (session 2)
Frequency: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Zacchiroli at Sat Jun 14 02:46:43 METDST 1997) Minor failures
Experiment lost first 6 hours. A very high atmosphere attenuation is present

Jodrell (Alastair Gunn at Mon Jun 16 11:42:29 METDST 1997) Minor failures
Observed through very heavy rain! FS crashed, missed scan 12 at 07:52 UT, ready for 08:01 UT.

Medicina ( at Mon Jun 16 14:38:23 METDST 1997) Failure
Probably lost first 9 hours due to error setting LO. Problems at 10:10, 18:26-18:34,21"39 and from 22:14 due to heavy rain.

Metsahovi (Kaj Wiik, Ari Mujunen at Tue Jun 17 15:02:49 METDST 1997) Success
Great weather, no apparent problems. This is the first EVN tape after our recorder was upgraded for thin tape. Two first minutes lost due to tracking computer problems. One suspicious head position at 97-164 07:58 pass/13,,auto,0.0,,191.8,,191.8. No formatter 1sec time jumps.

Effelsberg (dgraham at Wed Jun 18 10:10:19 METDST 1997) Minor failures
Some question about LO frequency for first 8 hours, probably OK. Awful rainy weather.

Onsala (John Conway at Wed Jun 18 14:01:00 METDST 1997) Minor failures
Setup instructions not followed, 1MHz filters connected to VCs 1,2,3,4, not sensible since we only have LCP from antenna feeding odd VCs and will miss the second frequency on VC5 for those shources which have it. Weather was fine however.

Noto (C.Trigilio at Thu Jun 19 12:08:10 METDST 1997) Success
Good weather, no problem.